What is the standard form of an ellipse whose endpoints of the major axis are at (-11,5) and (7,5) and endpoints of the minor axis are at (-2,9) and (-2,1)?
Answer by lwsshak3(11628)

What is the standard form of an ellipse whose endpoints of the major axis are at (-11,5) and (7,5) and endpoints of the minor axis are at (-2,9) and (-2,1)?
***
Standard form of equation for an ellipse with horizontal major axis:
(x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
For given ellipse:
center:(-2,5)
length of horizontal major axis=11+7=18=2a
a=9
a^2=81
length of minor axis=9-1=8=2b
b=1
b^2=1
Equation of given ellipse:
(x+2)^2/81+(y-5)^2=1